# BootSeer: Analyzing and Mitigating Initialization Bottlenecks in Large-Scale LLM Training

### v1

@@ Abstract
Large Language Models (LLMs) have become a cornerstone of modern AI, driving breakthroughs in natural language processing and expanding into multimodal jobs involving images, audio, and video. As with most computational software, it is important to distinguish between ordinary runtime performance and startup overhead. Prior research has focused on runtime performance: improving training efficiency and stability. This work focuses instead on the increasingly critical issue of startup overhead in training: the delay before training jobs begin execution. Startup overhead is particularly important in large, industrial-scale LLMs, where failures occur more frequently and multiple teams operate in iterative update-debug cycles. In one of our training clusters, more than 3.5% of GPU time is wasted due to startup overhead alone.
@@ 1 Introduction
In practice, large organizations often run many training jobs simultaneously, managed by sizable algorithm teams. These jobs frequently stop and restart due to debugging, system failures, node slowdowns, or iterative algorithm updates. For example, a recent training job in our platform on 4,864 NVIDIA H800 GPUs experienced 26 startups (10 of which are due to debugging) and 6 hot updates within 21 hours.
Figure 1: Breakdown of consumed GPU-server-hours into actual training (blue) and startup overhead (orange) in a cluster perspective. Aggregate data are collected from one of our training cluster over a single day.
Figure 1 shows a conservative estimate of the startup overhead in a training cluster, indicating that startup activities consume approximately 3.5% of total GPU server hours. This overhead negatively impacts AI development in three key ways: (1) it leads to inefficient GPU utilization; (2) it greatly impact the debug-and-resubmit cycle of development; and (3) excessively long initialization times can introduce unexpected stability issues.
@@ 2 Background
Section 2.1 discusses the process of training and releasing a new LLM in a large organization. This includes both startup and the actual training. The key insight is that training a large LLM model (often over a period of 6 months) is more akin to computational steering, rather than the naive concept of “submit a job and wait until it’s done”. Development unfolds in a sparse, tree-structured environment, in which multiple teams may concurrently fork the current main branch of the LLM, and test a modified algorithm. Only the most successful forks, typically those achieving significantly lower loss, are retained and further explored. Consequently, any training job that completes without bugs and achieves promising results becomes a candidate for the new main branch.
Section 2.2 then goes on to discuss the internals of the startup process. This includes: a Scheduler Phase; allocation of resources; loading an image; setting up the environment; and initializing the execution.
Finally, Section 2.3 points out where are the largest opportunities (“the low-hanging fruit”) for optimizing the startup phase in LLMs.
2.1 The process of training a new LLM Model for release in a large, industrial organization
@@
A typical introduction to building LLMs gives the impression that a single model is written in advance (perhaps in pyTorch), and the training phase executes until the loss of that model is sufficiently small. In this simplistic conception of building a model, one would imagine that if there is no failure of computer nodes, then the computation has one startup phase at the beginning, and the remaining computation is simply the training phase of an LLM computation as described in a typical introductory description. In this simplistic conception, the only reason to execute an additional startup phase is if there is a node failure, and the computation must revert to an earlier checkpoint.
The reality at industrial scale is quite different. Training jobs may be terminated for several reasons, including:
  • The model experienced a failure, either due to software or hardware issues.
  • The training was previously progressing well, but the loss suddenly increased for unknown reasons. In such cases, the job is terminated, reverted to a checkpoint where the loss was still decreasing, and restarted multiple times for debugging.
  • A team is experimenting with different post-training strategies. Multiple developer teams may apply various strategies to a base model. If a new approach fails to produce improvements, the job is terminated and the next idea is tried.
  • A higher-priority job has claimed the node resources, requiring the lower-priority job to be terminated.
With the above understanding, it is clear that the development of a new model might require on the order of a thousand different jobs each week. Thus, if we compare the development of a large model to High Performance Computing (HPC), the development of a large model is far from the concept of a “production job” in HPC, and is more akin to computational steering [29].
@@
2.2 The Startup Process
Startup can happen multiple time during a LLM training life-cycle, especially for those jobs using large-scale GPU cards. A Full Startup happens when a user: (i) first submits a training job; (ii) replaces a machine due to runtime failure/slowdowns; (iii) debugs the software. or (iv) changes a runtime configuration. Meanwhile, a Hot Update also causes overhead due to a partial startup (setting up an environment again and rerunning the model setup).
Figure 2 illustrates the simplified startup process of an LLM training job in our system. At a Scheduler Phase, the process appears similar to that of other distributed workloads. When a user submits a job, the scheduler takes over and allocates the required resources. Once the resources are provisioned, each worker node pulls the necessary container images, starts the container, and executes the setup or entrypoint script. Next, each machine initializes the model and begins the training process. However, certain stages may differ in detail compared to conventional workloads.
@@ 2.2 The Startup Process
Resource Allocation. The job scheduler (e.g., Kubernetes scheduler) is responsible for allocating resources to training jobs. Once sufficient resources are available — typically a set of physical GPU machines — the scheduler allocates them to the job and initiates the process of pulling the required container images.
Once worker nodes are ready, the startup process enters the Worker Phase:
Image Loading. In this stage, all worker nodes (GPU machines) begin pulling images concurrently from the container registry, and start the container. These images are typically much larger than those used in smaller jobs—ranging from 25 to 40 GB in our system. When the number of worker nodes is large, concurrent image loading can place significant strain on both the network bandwidth and the container registry service. If any sidecar containers are required (e.g., HDFS-FUSE), their images are also pulled during this stage.
Sometimes, a worker node may already have the training image cached locally, eliminating the need to pull it from the remote registry. However, since all machines must synchronize in the next stage, they are forced to wait for the slowest node to complete the image pull. As a result, a cache hit on a single machine does not significantly improve the overall efficiency of the image loading stage.
All experiments in this paper adopt the lazy-loading strategy instead of using standardized OCI images.
Environment Setup. In this stage, each worker node prepares its environment before running the actual training program (typically a Python program). First, it installs any dependencies that are not included in the container image. This is often necessary due to heterogeneous environments and the flexibility required by engineers. For instance, the version of certain dependencies may vary depending on the machine type, GPU type, operating system, or even geographic region. And some dependencies change often, so engineers prefer the flexibility to update them without rebuilding the container image.
@@
Model Initialization. In this stage, the user program begins execution, but effective model training has not yet started. This phase includes various steps that depend on the model architecture (e.g., Dense, MOE, RL, etc.). In general, it involves setting up the model according to the parallelization strategy, launching ranks (i.e., processes running on each GPU), loading model checkpoints, initializing RDMA connections, and other preparatory steps.
@@
2.3 Opportunities for Optimizing LLM Models
@@ 3.1 Impact of Startup Overhead on GPU Utilization and Training Efficiency
The first question we address in this section is: Why is startup overhead in LLM training important? We begin by quantifying how startup overhead contributes to underutilization of GPU resources and how this inefficiency scales with increasing job size.
@@ 3.2 Component-level Breakdown of Startup Overhead
Figure 5: Breakdown of node-level startup overhead across different initialization stages. Box plot whiskers extend to two standard deviations, in oorder to exclude outliers.
@@ 4.3 Environment Caching for Job-Level Dependency Installation
The Environment Setup phase, as introduced in Section 2.2, includes dependency installation and daemon initialization, which are critical for preparing the runtime environment. LauncControl installs package dependencies at this phase instead of placing all dependency packages in the Dockerfile because: (1) some dependency installations are decided based on runtime environment (e.g. GPU type, OS version); and (2) there are trade-off between flexibility and efficiency: some dependencies are upgraded very frequently and frequently building new container for new dependency versions would usually be time-consuming.
@@
The environment cache is generated by capturing file system changes during the initial environment setup. Figure 10 shows the process of creating and using the environment cache. In the first run of a job, Bootseer create the environment cache based on the difference of Target Directory before and after the Environment Setup phase on worker node 0, where the Target Directory is the dependnecy install path, e.g., the site-package directory. All files added or modified during the first environment setup are compressed and uploaded to HDFS through HDFS-FUSE.
@@ 4.4 Striped HDFS-FUSE for Efficient Checkpoint Resumption
Bootseer Optimizations: To overcome the I/O bottlenecks of baseline HDFS usage, Bootseer introduces a striped I/O mechanism for checkpoint access. Given the ample compute resources available on training nodes during the startup stage, the worker node have redundant resources to donwload data chunks in parallel. Figure 11 illustrates the layout of the striped HDFS-FUSE system.
@@ 5 Evaluation
In this section, we conduct end-to-end experiments to evaluate the training startup optimizations provided by Bootseer, in comparison to a baseline system. To evaluate the effectiveness of individual optimizations in Bootseer, we conducted a series of micro-benchmarks targeting key components described in Section 4. Therefore, unlike the previous sections, these experiments are concerned only with the startup phase, and do not represent production jobs.
@@ 5.1 Experiment Setup
Training Workload. We use large-scale MOE training jobs with significant checkpoint and container sizes to evaluate startup overhead under realistic, heavy-load conditions Each job trains an 8-layer Mixture-of-Experts (MOE) model with 128 experts per layer. The model is configured with 2-way pipeline parallelism, and the checkpoint size is 413 GB. We scale the training across 16, 32, 48, 64, and 128 GPU cards, corresponding to data parallelism degrees of 1, 2, 3, 4, and 8, respectively. The container image size for this job is 28.62 GB.
@@ 5.2 End-to-End Startup Optimization
To answer RQ1, we conduct an end-to-end experiment to evaluate the optimization performance of Bootseer and compare it with the baseline. Both frameworks are used to start the MOE training job described in Section 5.1. All results are averaged over three independent experiments. Before each run, we clear the local image cache to eliminate any potential impact.
To explain the performance gap, we detail the startup mechanisms used in both the baseline and Bootseer configurations
Baseline. The baseline follows our original training startup process:
  • The container image is downloaded using a lazy-loading mechanism, with peer-to-peer sharing enabled across all worker nodes.
  • Dependencies are installed during the environment setup phase.
  • Checkpoint files are retrieved from HDFS and loaded into memory to resume training.
Bootseer. The Bootseer setup includes the following components:
  • Record-and-prefetch for container images: This feature captures hot data blocks based on access patterns from previous runs using the same container image. A 2-minute window is used to record hot blocks, while cold blocks are prefetched in the background using 8 threads.
  • Environment cache: Environment caching is enabled, with cache files generated from previous executions of the same task. The compressed cache size is 270 MB and is stored in HDFS. These files are fetched and restored during the environment setup phase.
  • Checkpoint resumption: Checkpoint files are mounted to a local directory using striped HDFS-FUSE. This mounting occurs during the environment setup phase. HDFS-FUSE is supported by an auxiliary container, which must also be pulled during the image loading stage.
@@
Bootseer consistently reduces end-to-end startup time by 2 times compared to the baseline, and this benefit holds across different job scales. Figure 12 presents the startup overhead of both frameworks with respect to the scale of the training jobs. The startup overhead increases with the size of the training cluster, especially when scaling from 64 to 128 GPUs. This trend aligns with the findings discussed in Section 3.
@@ 5.3 Improvement Breakdown
For the Image Loading stage, Bootseer outperforms the baseline by a factor of 4 to 10 as the job scale increases. This improvement is due to the inefficiency of the lazy-loading setup used in the baseline: when a container accesses a data block that has not yet been loaded onto the worker node, it must fetch the block from a remote cache or another machine. As the job scale increases, these cache misses place additional pressure on the network within the training cluster. In contrast, Bootseer’s record-and-prefetch mechanism avoids remote data access, keeping duration stable regardless of job scale.
For the Environment Setup stage, Bootseer reduces latency by a factor of 2 by skipping most dependency installation commands. We also observe a noticeable increase in duration when scaling from 64 to 128 GPUs. This is because all machines must establish connections and synchronize with each other, and as the job scale increases, the associated overhead also grows. Bootseer does not include specific optimizations for this aspect.
For the Model Initialization stage, Bootseer reduces overhead by approximately 1.6× through faster checkpoint resumption. Compared to other stages, the duration is more stable and does not show a significant increase as the job scale grows. This may be because the current job scale is not large enough to trigger a bottleneck in our HDFS.
@@ 5.4 Troubleshooting Startup Stragglers
To address RQ3, we analyze how our environment cache helps mitigate stragglers in large-scale training. As repeatedly emphasized throughout this paper, access to remote resources is a major cause of stragglers during training startup. This observation motivated the design of a job-level environment cache, aimed at reducing startup overhead while preserving the necessary flexibility.
Figure 14: Distribution of execution times (duration) for the dependency installation script across all nodes in the 128-GPU experiment. The whiskers of the box extend to the minimum and maximum duration values.
Bootseer significantly reduces variability in dependency installation duration, effectively eliminating startup stragglers. Figure 14 presents the distribution of execution times for the dependency installation script that runs during the Environment Setup stage. We collected the execution times from each worker node in the 128-GPU training job and visualized the results using a boxplot. The results in this figure demonstrate that Bootseer not only reduces package installation overhead, but also eliminates stragglers.
The value of environment caching becomes more significant as job scale increases, where straggler effects grow more severe. Although the straggler effect in this experiment is based on just a few hundred GPUs, the results in Section 3.3 indicate that the effect becomes more pronounced as the job scales to hundreds or thousands of GPU servers. In such large-scale settings, the environment cache becomes even more valuable.

### v2

@@ Abstract
Large Language Models (LLMs) have become a cornerstone of modern AI, driving breakthroughs in natural language processing and expanding into multimodal jobs involving images, audio, and video. As with most computational software, it is important to distinguish between ordinary runtime performance and startup overhead. Prior research has focused on runtime performance: improving training efficiency and stability. This work focuses instead on the increasingly critical issue of startup overhead in training: the delay before training jobs begin execution. Startup overhead is particularly important in large, industrial-scale LLMs, where failures occur more frequently and multiple teams operate in iterative update-debug cycles. In one of our production training clusters, startup overhead alone wastes over 4,700 GPU-hours in a single day.
@@ 1 Introduction
In practice, large organizations often run many training jobs simultaneously, managed by sizable algorithm teams. These jobs frequently stop and restart due to debugging, system failures, node slowdowns, or iterative algorithm updates. For example, a recent training job on our platform running on 4,864 NVIDIA H800 GPUs experienced 26 full startups (10 due to debugging) and 6 hot updates within a 21-hour window.
Although each startup is transient, their cumulative impact is substantial. Slow startup phases significantly degrade development efficiency by lengthening the debug–modify–resubmit cycle. Instead of iterating rapidly on model or system changes, engineers must repeatedly wait for long initialization phases before observing any training progress. This delay slows experimentation, reduces effective developer throughput, and exacerbates coordination overhead across large teams. Moreover, prolonged and repeated initialization increases exposure to transient system issues, further disrupting the development process.
@@ 2 Background
Section 2.1 discusses the process of training and releasing a new LLM in a large organization. Section 2.2 then goes on to discuss the internals of the startup process. This includes: a Scheduler Phase; allocation of resources; loading an image; setting up the environment; and initializing the execution. Finally, Section 2.3 points out where are the largest opportunities (“the low-hanging fruit”) for optimizing the startup phase in LLMs.
2.1 LLM Training in Production
@@
A simplistic conception of building an LLM model posits that if there is no failure of computer nodes, then the computation has just one startup phase at the beginning. The remaining computation is simply to continue training until the loss for the LLM model is sufficiently small. In this simplistic conception, the only reason to execute an additional startup phase is if there is a node failure, and the computation must revert to an earlier checkpoint.
At industrial scale, the reality is quite different. Training jobs may be terminated for several reasons, including:
  • The model experienced a software or hardware failure.
  • The training was previously progressing well, but the loss suddenly increased for unknown reasons. In such cases, the job is terminated. The model is then reverted to a checkpoint where the loss was still decreasing, and restarted multiple times for debugging.
  • A team is experimenting with different post-training strategies. Multiple developer teams may apply various strategies to a base model. If a new approach fails to quickly produce improvements, the job is terminated and the next idea is tried.
  • A higher-priority job has claimed the node resources, requiring the lower-priority job to be aborted.
@@
2.2 Types of Startups
Startups happen multiple times and for multiple reasons during an LLM training life-cycle, especially for those jobs using large-scale GPU cards. A Full Startup happens when a user: (i) first submits a training job; (ii) replaces a machine due to runtime failure/slowdowns; (iii) debugs the software; or (iv) changes a runtime configuration. Meanwhile, a Hot Update also causes overhead due to a Partial Startup (setting up an environment again and rerunning the model setup).
Figure 1 illustrates the simplified startup process of an LLM training job. In the Scheduler Phase, the cluster operation appears similar to that of other distributed workloads. When a user submits a job, the scheduler provisions the required resources. Each worker node then “pulls” the necessary container images, starts the container, and executes the setup or entrypoint script. Next, each machine initializes the model and begins the training process.
@@ 2.2 Types of Startups
Resource Allocation. The job scheduler (e.g., Kubernetes [8, 28]) is responsible for allocating resources for training jobs. Once sufficient resources are available — typically a set of physical GPU machines — the scheduler allocates them to the job and initiates the process of “pulling” the required container images.
The startup process then enters the Worker Phase:
Image Loading. In this stage, all worker nodes (GPU machines) begin pulling images concurrently from the container registry, and then start the containers. These images are typically much larger than those used in smaller jobs — ranging from 25 to 40 GB in our system. When the number of worker nodes is large, concurrent image loading places significant strain on both the network bandwidth and the container registry service. If any sidecar containers are required (e.g., HDFS-FUSE), their images are also pulled during this stage.
All machines must then wait for the slowest node to complete the image pull, in order to synchronize. Sometimes, a worker node may already have the training image cached locally, eliminating the need to pull it from the remote registry. Nevertheless, a cache hit on a single machine does not necessarily improve the overall efficiency at this stage. All experiments in this paper adopt the lazy-loading strategy instead of using standardized OCI images.
Environment Setup. In this stage, each worker node prepares its environment before running the actual training program (typically a Python program). First, it installs any dependencies that are not included in the container image. This is often necessary due to heterogeneous environments and the flexibility required by engineers. For instance, the version of certain dependencies may vary depending on the machine type, GPU type, operating system, or even geographic region. And some dependencies change often, because engineers frequently update them, without rebuilding the container image.
@@
Model Initialization. In this stage, the user program begins execution, but effective model training has not yet started. In this stage, several steps depend on the model architecture (e.g., Dense, MoE, RL, etc.). This involves setting up the model according to the parallelization strategy, launching ranks (i.e., processes running on each GPU), loading model checkpoints, initializing RDMA connections, among other steps.
Startup Cost Breakdown in Practice. To illustrate the cost of the above startup phases in real production settings, we report measurements from two recent industrial LLM training jobs. All timings are measured at the job level.
The first job trains a pure language MoE base model on 1,440 machines (11,520 NVIDIA GPUs). The initial worker process launch takes 21 seconds. Model initialization (including imports, model graph construction, parallelism setup, and NCCL initialization) takes 41 seconds, within which NCCL initialization alone accounts for 6.7 seconds. Resuming from a checkpoint takes 3 minutes and 13 seconds, and data loader setup takes 18.5 seconds.
The second job trains a multimodal model on 384 machines, with a substantially larger checkpoint that includes vision model parameters. In this case, the initial worker process launch takes 10 seconds, model initialization takes 22 seconds, checkpoint resume takes 10 minutes and 12 seconds, and data loader setup takes 5 minutes and 17 seconds. These results show that startup overheads can vary significantly across models and configurations, and that checkpoint and data loading phases can dominate the end-to-end startup cost.
@@
2.3 Opportunities for Optimizing Training Startup
@@ 3.1 Impact of Startup Overhead on GPU Utilization and Training Efficiency
The first question we address in this section is: Why is startup overhead in LLM training important? We begin by quantifying how startup overhead contributes to the underutilization of GPU resources and how this inefficiency scales with increasing job size.
@@ 3.2 Component-level Breakdown of Startup Overhead
Figure 4: Breakdown of node-level startup overhead across different initialization stages. Box plot whiskers extend to two standard deviations, in order to exclude outliers.
@@ 4.3 Environment Caching for Job-Level Dependency Installation
The Environment Setup phase, as introduced in Section 2.2, includes dependency installation and daemon initialization, which are critical for preparing the runtime environment. BootSeer installs package dependencies at this phase instead of placing all dependency packages in the Dockerfile because: (1) some dependency installations are decided based on runtime environment (e.g. GPU type, OS version); and (2) there are trade-off between flexibility and efficiency: some dependencies are upgraded very frequently and frequently building new container for new dependency versions would usually be time-consuming.
@@
The environment cache is generated by capturing file system changes during the initial environment setup. Figure 9 shows the process of creating and using the environment cache. In the first run of a job, BootSeer create the environment cache based on the difference of Target Directory before and after the Environment Setup phase on worker node 0, where the Target Directory is the dependency install path, e.g., the site-package directory. All files added or modified during the first environment setup are compressed and uploaded to HDFS through HDFS-FUSE.
@@ 4.4 Striped HDFS-FUSE for Efficient Checkpoint Resumption
BootSeer Optimizations: To overcome the I/O bottlenecks of baseline HDFS usage, BootSeer introduces a striped I/O mechanism for checkpoint access. Given the ample compute resources available on training nodes during the startup stage, the worker node have redundant resources to download data chunks in parallel. Figure 10 illustrates the layout of the striped HDFS-FUSE system.
@@ 5.1 Experiment Setup
Training Workload. We use large-scale MoE training jobs with significant checkpoint and container sizes to evaluate startup overhead under realistic, heavy-load conditions Each job trains an 8-layer Mixture-of-Experts (MoE) model with 128 experts per layer. The model is configured with 2-way pipeline parallelism, and the checkpoint size is 413 GB. We scale the training across 16, 32, 48, 64, 128, 512 and 1024 GPU cards, corresponding to data parallelism degrees of 1, 2, 3, 4, 8, 16 and 32, respectively. The container image size for this job is 28.62 GB. Note that the evaluation task we selected is a training job typically used at large scales (1000+ GPUs), and therefore does not align with the statistical characteristics described in Section 3.
@@ 5.2 End-to-End Startup Optimization
To answer RQ1, we conduct an end-to-end experiment to evaluate the optimization performance of BootSeer compared to the baseline. Both frameworks are used to start the MoE training job described in Section 5.1. All results are averaged over three independent experiments. Before each run, we clear the local image cache to eliminate potential impacts.
To explain the performance gap, we detail the startup mechanisms used in both the baseline and BootSeer.
Baseline. The baseline follows our original training startup process: (1) the container image is downloaded using a lazy-loading mechanism with peer-to-peer sharing across all worker nodes, (2) dependencies are installed during startup, and (3) checkpoint files are retrieved from HDFS and loaded into memory to resume training.
BootSeer. The BootSeer setup includes the following components: (1) Record-and-prefetch for container images — Hot data blocks are identified based on access patterns from previous runs using the same container image. A two-minute window is used to capture hot blocks, while cold blocks are prefetched in the background using eight threads. (2) Environment cache — Environment caching is enabled, with cache files generated from previous executions of the same task. The compressed cache (270 MB) is stored in HDFS and restored during environment setup. (3) Checkpoint resumption — Checkpoint files are mounted to a local directory using striped HDFS-FUSE during the setup phase. HDFS-FUSE is supported by an auxiliary container that must also be pulled during image loading.
@@
BootSeer consistently reduces end-to-end startup time by 2 times compared to the baseline, and this benefit holds across different job scales. Figure 11 presents the startup overhead of both frameworks with respect to the scale of the training jobs. The startup overhead increases with the size of the training cluster, especially when scaling from 64 to 1024 GPUs. This trend aligns with the findings discussed in Section 3.
@@ 5.3 Improvement Breakdown
Figure 13: Distribution of execution times for the dependency installation script (a component of the Environment Setup stage) across varying job scales.
Image Loading. BootSeer outperforms the baseline by a factor of 4× to 7× as job scale increases. This improvement stems from inefficiencies in the baseline’s lazy-loading mechanism: when a container accesses a data block that has not yet been loaded onto the local node, it must retrieve the block from a remote cache or another machine. As the number of GPUs increases, such cache misses place additional strain on the cluster’s network. In contrast, BootSeer’s record-and-prefetch strategy eliminates remote data access, maintaining stable loading times regardless of scale.
Environment Setup. BootSeer reduces latency by 20% to 50% by bypassing most dependency installation commands. However, as the job scale increases from 128 to 1024 GPUs, we observe a significant rise in duration and diminishing performance gains. This trend is attributed to the growing overhead of establishing and synchronizing connections across all nodes, a process that BootSeer does not explicitly optimize.
Model Initialization. BootSeer reduces overhead by up to 2.6× through more efficient checkpoint resumption. Among all stages, Model Initialization shows the most stable performance, with no substantial increase in duration as job scale grows. This stability may be due to the current job sizes not being large enough to cause bottlenecks in our HDFS.
@@ 5.4 Troubleshooting Startup Stragglers
To address RQ3, we analyze how our environment cache helps mitigate stragglers in large-scale training. As repeatedly emphasized throughout this paper, access to remote resources is a major cause of stragglers during training startup. This observation motivates the design of a job-level environment cache, aimed at reducing startup overhead while preserving the necessary flexibility.
BootSeer significantly reduces variability in dependency installation times, effectively eliminating startup stragglers. Figure 13 shows the distribution of execution times for the dependency installation script executed during the Environment Setup stage. Execution times were collected from each worker node across different job scales. The results demonstrate that BootSeer not only reduces package installation overhead, but also eliminates stragglers.
The value of environment caching becomes more significant as job scale increases, where straggler effects grow more severe. Although the straggler effect in this experiment is based on just a few hundred GPUs, the results in Section 3.3 indicate that the effect becomes more pronounced as the job scales to hundreds or thousands of GPU servers.